40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2650     Entry time: Wed Feb 17 13:25:52 2021
Author: aaron 
Type: DailyProgress 
Category: PSOMA 
Subject: electronics management 

enter Wed Feb 17 13:25:50 2021

I'm clearing electronics from under the PSOMA table, and temporarily storing them on the lH2 dewar's base. It's quite dusty under there, and I wiped down what was below. I've added the wire shelf we want to order to the PSOMA hardware inventory ('other'), along with 80-20 components to make a tabletop "soundproof" enclosure.

While under the table, I noticed the N2 line (for floating the table) is touching a strut attached to the floated table, creating a mechanical short. I've added zipties to avoid this. Also mounted the power strips to the table legs' struts, and labelled power cables as I went.

The MA-40 HEPA filter unit indicates it's time for a filter replacement. I've contacted Calum, who might have a supply of spare filters for these units. Perhaps it's also time to check the FFU on cryo cavs table.

Began but did not finish rerouting cable on the tabletop.
 

exit Wed Feb 17 20:18:39 2021

ELOG V3.1.3-